the sum of 3 consecutive integers is 60 how do you solve that?
Answer by stanbon(75887)

the sum of 3 consecutive integers is 60 how do you solve that?
---
1st: x-1
2nd: x
3rd: x+1
-----------
Equation:
sum = 60
3x = 60
x = 20
----
1st = x-1 = 19
2nd = x = 20
3rd: x+1 = 21
==================
Cheers,
Stan H